OpenStudy (anonymous):

what is the equation of a line passing through (-8, -3) and having a slope of -4? can someone walk me through it

OpenStudy (noseboy908):

For this problem you're gonna have to use point-slope form. Do you know what that is?

OpenStudy (anonymous):

yea its y-y1=m(x-x1)

OpenStudy (anonymous):

y1 would be -3? and x1 be -8

OpenStudy (noseboy908):

yep

OpenStudy (noseboy908):

and the slope would be?

OpenStudy (anonymous):

okqy slope of -4 but they want this in y=nx+b

OpenStudy (noseboy908):

*m

OpenStudy (anonymous):

so how do i do that ?

OpenStudy (noseboy908):

so fill in what you know into point-slope

OpenStudy (anonymous):

that makes no sense to me how do i go from point slope to y=mx+b

OpenStudy (noseboy908):

just fill it in, and we'll go from there. You cannot just put in a y-int, you don't have it. So you have to solve for it

OpenStudy (anonymous):

y=-4x+ how do i solve to get b

OpenStudy (noseboy908):

no, fill in POINT-SLOPE form, not slope-int

OpenStudy (anonymous):

y-3=-4x-3

OpenStudy (noseboy908):

what happened to the -8?

OpenStudy (anonymous):

y--3=4x+8

OpenStudy (noseboy908):

no, one step at a time

OpenStudy (noseboy908):

rewrite it with just the variables filled in

OpenStudy (anonymous):

okay y--3=-4x-8

OpenStudy (noseboy908):

no, it's "m(x-x1)" not " mx-x1"

OpenStudy (anonymous):

y--3=-4(x-8)

OpenStudy (noseboy908):

correct, (just rewrite "y--3" as "y+3). so now distribute the -4

OpenStudy (anonymous):

y+3=-4x+32

OpenStudy (noseboy908):

yeah! now isolate the Y.

OpenStudy (anonymous):

y=-4x+29

OpenStudy (noseboy908):

correct!

OpenStudy (anonymous):

but there is no -29 as an answer choice

OpenStudy (noseboy908):

although, looking back, you could have left it in point-slope form

OpenStudy (noseboy908):

It doesn't ask for a particular form

OpenStudy (noseboy908):

so, you're done!

OpenStudy (anonymous):

= −4x + 20B. y = 4x + 20C. y = −4x – 35D. y = 4x – 35E. y = -4x + 4

OpenStudy (anonymous):

here are the answer choices

OpenStudy (noseboy908):

rechecking, you did one part wrong. redistribute the -4

OpenStudy (anonymous):

ok so what did i do

OpenStudy (noseboy908):

redistribute the -4. see what you get

OpenStudy (anonymous):

im still getting -32

OpenStudy (noseboy908):

that's right, but what else do you have? rewrite your formula that you get right after you distribute

OpenStudy (anonymous):

or do i distrubute to the other side to

OpenStudy (noseboy908):

just rewrite your equation

OpenStudy (anonymous):

y=4x+20

OpenStudy (noseboy908):

no, the equation as right after you distributed the -4

OpenStudy (anonymous):

y--3=-4x-32

OpenStudy (noseboy908):

again, make sure to rewrite the "y--3" as "y+3". but next, to isolate the y, what do you do?

OpenStudy (anonymous):

add 3 to both sides

OpenStudy (noseboy908):

no, this is why it's important to make the 2 negatives a positive. when you have "y+3..." you have to SUBTRACT the 3 to isolate y

OpenStudy (anonymous):

so it would be y=-4x-35?

OpenStudy (noseboy908):

yes!

OpenStudy (anonymous):

thank you !

OpenStudy (noseboy908):

welcome! make sure to hit "best response"! it helps!